OpenStudy (anonymous):

g+1/g-4 divided by 7g+7/g+1 Divide and simplify

OpenStudy (anonymous):

To find the simplified form of a complex fraction, multiply by the form of 1=a/a where a is the LCD of all of the inner fractions.

OpenStudy (lalaly):

\[\frac{g+1}{g-4} \times \frac{g+1}{ 7g+7}\]

OpenStudy (anonymous):

\[\frac{\frac{1}{p}}{\frac{r}{q}} \cdot \frac{pq}{pq} = \frac{q}{pr}\]

OpenStudy (anonymous):

\[\large \qquad \frac{\frac{g+1}{g-4}}{\frac{7g+7}{g+1}}\]\[=\large \frac{\frac{g+1}{g-4}}{\frac{7\cancel{(g+1)}}{\cancel{g+1}}}\]\[=\large \frac{\frac{g+1}{g-4}}{7} \cdot \frac{g-4}{g-4}\qquad; g \ne -1\]\[=\large\frac{g+1}{7(g-4)}\qquad \qquad; g \ne -1\]